1. A company had 80 employees whose salaries are summarized in the frequency distribution below.

Salary ($) Employees 5,001-10,000 10 10,001-15,000 16 15,001-20,000 18 20,001-25,000 11 25,001-30,000 25  

Find the mean salary.

a. $20,969.30

b. $17,156.70

c. $17,500

d. $19,063.00
Answer #1
Salary Class mark (x) Frequency(f) x*f
5001-10000 7500 10 75000
10001-15000 12500 16 200000
15001-20000 17500 18 315000
20001-25000 22500 11 247500
25001-30000 27500 25 687500
Total 80 1525000

mean = \frac{\sum x*f}{\sum f} = \frac{1525000}{80} = \$ 19,063

Answer d. $19,063.00

​Please give thumbs up if you like the answer.